# Standardized Constraints on the Shadow Radius and the Instability of Scalar, Electromagnetic, $p$-Form, and Gravitational Perturbations of High-Dimensional Spherically Symmetric Black Holes in Einstein-power-Yang-Mills-Gauss-Bonnet Gravity

### v4

@@ Abstract
The space-time geometry under investigation is chosen to be a high-dimensional, static, spherically symmetric solution in an asymptotically flat background within the Einstein-power-Yang-Mills-Gauss-Bonnet (EPYMGB) gravity. To address the limitations of previous shadow constraints, we construct a standardized framework based on the Schwarzschild-Tangherlini metric to constrain the characteristic parameters of high-dimensional black holes by leveraging observational shadow data. Additionally, we provide a rigorous derivation of the shadow radius formula for a general high-dimensional spherically symmetric black hole. Subsequently, we systematically and comprehensively present the equations of motion and master variables governing spin-0, spin-1, $p$-form, and spin-2 perturbations in high-dimensional static spherically symmetric flat space-time. Our analysis reveals that the Yang-Mills magnetic charge $\mathcal{Q}$ and the power $q$ have a negligible impact on both the shadow radius and perturbations of the black hole when compared to the Gauss-Bonnet coupling constant ${\alpha}_{2}$ in various dimensions. Hence, the physical signatures of the parameters $\mathcal{Q}$ and $q$ in the black hole environment remain undetectable through either perturbation analysis or shadow observations. Cross-validation of the allowable range of ${\alpha}_{2}$ derived from the high-dimensional constraint on shadow radius and the dynamical stability analysis of gravitational perturbations demonstrates excellent agreement between these independent approaches. The conclusion of this cross-analysis further substantiate the accuracy of the high-dimensional shadow constraint formula proposed in this work, and we argue that this formula may serve as a universal tool for constraining the characteristic parameters of other high-dimensional spherically symmetric black hole solutions.
@@ II.1 The action and the field equations
$\mathcal{S}=\frac{1}{2}\int_{\mathcal{M}}\mathrm{d}^{n}x\sqrt{-g}\left[\mathcal{R}-\frac{(n-2)(n-1)}{3}\Lambda+{\alpha}_{2}\mathcal{L}_{\text{GB}}-\mathcal{F}^{q}\right],$ (1)
where $8\pi\mathscr{G}=1$, the parameter ${\alpha}_{2}$ is the Gauss-Bonnet coupling constant, the Gauss-Bonnet Lagrangian $\mathcal{L}_{\text{GB}}$ is given as
$\mathcal{L}_{\text{GB}}=\mathcal{R}_{\gamma\zeta\varpi\delta}\mathcal{R}^{\gamma\zeta\varpi\delta}-4\mathcal{R}_{\gamma\zeta}\mathcal{R}^{\gamma\zeta}+\mathcal{R}^{2},$ (2)
and the Yang-Mills invariant $\mathcal{F}$ is represented as
@@
where $\mathbf{A}^{(a)}$ or $A_{\zeta}{}^{(a)}$ are the Yang-Mills potentials of the gauge group SO(n1)SO(n-1), $\sigma$ is the coupling constant and $C_{(b)(c)}^{(a)}$ are interpreted as the structure constants of $\frac{(n-1)(n-2)}{2}$-parameter Lie group. The field equation of the space-time metric $g_{\mu\nu}$ and the gauge potential $A_{\zeta}{}^{(a)}$ after the variation of action can be explicitly expressed as
@@ II.2 The high-dimensional black hole solution
The metric for high-dimensional ($n\geq 5$) static spherically symmetric is ansatz
@@
where $\mathfrak{h}_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ represents the line element of a (n2)(n-2)-dimensional hypersurface with the volume $\Sigma_{(n-2)}$. The solution for a EGBPYM black hole in asymptotically flat space-time can be written as follows:
@@
This metric is the generalization of Schwarzschild space-time in higher dimensions cases, which is proved to be Ricci-flat. When considering an uncharged black hole, it can be observed that the EGBPYM solution (17) can be transformed into the Schwarzschild-Tangherlini solution (18) by applying the limit condition ${\alpha}_{2}\rightarrow 0$. When no limit condition is imposed on ${\alpha}_{2}$, the uncharged EGBPYM solution (17) is the EGB solution Cai (2002) in asymptotically flat space-time, which is
@@
When the limit ${\alpha}_{2}\rightarrow 0$ is set and $q=1$ is selected, the EGBPYM solution (17) will be reduced to the EYM solution Guo and Miao (2020) as follows:
@@
In order for the EGBPYM black hole solution to have at least one event horizon, that is, to ensure that the black hole does not expose a naked singularity, the values of the parameters ${\alpha}_{2}$ and $\mathcal{Q}$ are restricted to a fixed region, as shown in Figures 2 and 3. It is worth noting that in the process of numerical calculation of this fixed region, the specific conditions of the metric (15) should also be considered as follows:
@@ III.2 Standardization of shadow constraints
$\theta_{\mathrm{g}}=\frac{\mathsf{G}M}{c^{2}D},$ (73)
where $\mathsf{G}$ represents the gravitational constant term, $M$ is the black hole mass inferred from the EHT and $D$ is the observed distance between the black hole and the observer. Next, the fractional deviation of the angular gravitational radius is introduced as follows:
@@
where $\theta_{\mathrm{g}}$ and $\theta_{\mathrm{dyn}}$ are used to represent the real measurement and the inference of the stellar-dynamics of the angular gravitational radius, respectively. In other words, $\theta_{\mathrm{g}}$ and $\theta_{\mathrm{dyn}}$ correspond to the measured value and estimated value, respectively. Thus, $\theta_{\mathrm{dyn}}$ can be written as
$\theta_{\mathrm{dyn}}=\frac{\mathsf{G}M_{\mathrm{dyn}}}{c^{2}D_{\mathrm{dyn}}}.$ (75)
@@ III.2.2 High-dimensional space-time
Next, we define the function $\mathfrak{C}(n)$ as
@@
$=\mathfrak{C}(n)\cdot\left(\mathscr{G}M_{\mathrm{dyn}}\right)^{\frac{1}{n-3}-1}\cdot\left(\frac{\mathscr{G}M_{\mathrm{dyn}}}{D_{\mathrm{dyn}}}\right)$
$=\mathfrak{C}(n)\cdot\left(\mathscr{G}M_{\mathrm{dyn}}\right)^{\frac{1}{n-3}-1}\cdot\theta_{\mathrm{dyn}}.$
@@
$\delta=\frac{\left[\frac{\theta_{\mathrm{sh}}}{\mathfrak{C}(n)\cdot\left(\mathscr{G}M_{\mathrm{dyn}}\right)^{\frac{1}{n-3}-1}}\right]}{\theta_{\mathrm{g}}}-1,$ (88)
@@
$\theta_{\mathrm{sh}}=\left[\mathfrak{C}(n)\cdot\left(\mathscr{G}M_{\mathrm{dyn}}\right)^{\frac{1}{n-3}-1}\right]\left(\delta+1\right)\theta_{\mathrm{g}}.$ (89)
@@
$\theta_{\mathrm{sh}}=\frac{r_{\mathrm{sh}}}{D}=\frac{r_{\mathrm{sh}}}{\left(\frac{\mathscr{G}M}{\theta_{\mathrm{g}}}\right)}=\left[\mathfrak{C}(n)\cdot\left(\mathscr{G}M_{\mathrm{dyn}}\right)^{\frac{1}{n-3}-1}\right]\left(\delta+1\right)\theta_{\mathrm{g}},$ (90)
@@
$r_{\mathrm{sh}}=\left[\mathfrak{C}(n)\cdot\left(\mathscr{G}\right)^{\frac{1}{n-3}}\cdot\left(M_{\mathrm{dyn}}\right)^{\frac{1}{n-3}-1}\right]\left(\delta+1\right)M.$ (91)
It should be noted here that $\mathscr{G}$ and $M_{\mathrm{dyn}}$ depend on the dimension $n$. We can refer to the extra-dimensional compactification volume $\mathscr{V}_{(n-4)}$ to relate the high-dimensional gravitational constant $\mathscr{G}$ to the four-dimensional gravitational constant $G$, and the relationship is defined in terms of
$G=\frac{\mathscr{G}}{\mathscr{V}_{(n-4)}}.$ (92)
If the extra dimension is a flat torus with $n$ dimensions, then $\mathscr{V}_{(n-4)}=\left(2\pi\mathscr{R}\right)^{n-4}$. Thus, the high-dimensional gravitational constant $\mathscr{G}$ can be expressed as
$\mathscr{G}=G\left(2\pi\mathscr{R}\right)^{n-4},$ (93)
where $\mathscr{R}$ represents the extra-dimensional compactification radius for each dimension $n$. Accordingly, the shadow radius $r_{\mathrm{sh}}$ can be rewritten as
$r_{\mathrm{sh}}=\left[\mathfrak{C}(n)\cdot\left(G\right)^{\frac{1}{n-3}}\cdot\left(\frac{M_{\mathrm{dyn}}}{2\pi\mathscr{R}}\right)^{\frac{1}{n-3}-1}\right]\left(\delta+1\right)M.$ (94)
If we consider stipulating $M_{\mathrm{dyn}}=2\pi\mathscr{R}$ for each dimension $n$, the above formula can be simplified to
$r_{\mathrm{sh}}=\left[\mathfrak{C}(n)\cdot\left(G\right)^{\frac{1}{n-3}}\right]\left(\delta+1\right)M.$ (95)
Since the Gauss-Bonnet term serves merely as a higher-order curvature correction, the four-dimensional gravitational constant $G$ after compactification can still be expressed as the ratio of the high-dimensional gravitational constant $\mathscr{G}$ to the extra-dimensional compactification volume $\mathscr{V}_{(n-4)}$. As defined above, it follows clearly that $M_{\mathrm{dyn}}\equiv\mathsf{M}$. Therefore, the $\mathscr{G}\mathsf{M}$” term in the metric (17) can also be reformulated as $G\left(2\pi\mathscr{R}\right)^{n-3}$” in compliance with the stipulation $\mathsf{M}=2\pi\mathscr{R}$. In practical calculations, in order to obtain the value of $R_{s}$ (i.e., Eq. (71)) corresponding to the metric (17), we will further specify $\mathsf{M}=2\pi\mathscr{R}=1$ for convenience. Note that this specification implies a fixed value selected when calculating a certain dimension $n$, rather than indicating that it holds for different dimensions. When we do not consider calculations between different dimensions, the above specification is allowed.
@@
The shadow radius constraint boundaries $r_{\mathrm{sh}}$ in high-dimensional space-time exhibit explicit dimensional dependence, with distinct parameter bounds emerging for each space-time dimension $n$. Additionally, an intriguing phenomenon is that for the high-dimensional constraint equation, the contribution of $G$ to $r_{\mathrm{sh}}$ is affected by the dimension $n$ when $G\neq 1$.
Therefore, for the sake of standardization of shadow radius constraints, when we systematically constrain the characteristic parameters in high-dimensional black hole solutions, Eq. (95) should be used.
@@ V.2 High-dimensional constraint results of shadow radius
Firstly, we apply the high-dimensional shadow radius formula (71) constructed in this paper to investigate the EPYMGB black hole solutions, and we present the corresponding numerical result of $R_{s}$ within the allowed parameter space $\left({\alpha}_{2},\mathcal{Q}\right)$. Subsequently, we utilize the high-dimensional shadow constraint formula (95) to provide constraints on the previously obtained values of $R_{s}$ based on observational data from the $\mathrm{M87}^{\star}$ and $\mathrm{Sgr~A}^{\star}$, respectively. The results are shown in Figures 4, 5, 6 and 7. Among these results, we set the ADM mass of the black hole and the four-dimensional gravitational constant as $\mathsf{M}=M=1$ and G=1{G}=1, respectively.

### v5

@@ Abstract
The space-time geometry under investigation is chosen to be a high-dimensional, static, spherically symmetric solution in an asymptotically flat background within the Einstein-power-Yang-Mills-Gauss-Bonnet (EPYMGB) gravity. To address the limitations of previous shadow constraints, we construct a standardized framework based on the Schwarzschild-Tangherlini metric to constrain the characteristic parameters of high-dimensional black holes by leveraging observational shadow data. Additionally, we provide a rigorous derivation of the shadow radius formula for a general high-dimensional spherically symmetric black hole. Subsequently, we systematically and comprehensively present the equations of motion and master variables governing spin-0, spin-1, $p$-form, and spin-2 perturbations in high-dimensional static spherically symmetric flat space-time. Our analysis reveals that the Yang-Mills magnetic charge $\mathcal{Q}$ and the power $q$ have a negligible impact on both the shadow radius and perturbations of the black hole when compared to the Gauss-Bonnet coupling constant ${\alpha}_{2}$ in various dimensions. Hence, the physical signatures of the parameters $\mathcal{Q}$ and $q$ in the black hole environment remain undetectable through either perturbation analysis or shadow observations. Cross-validation of the allowable range of ${\alpha}_{2}$ derived from the high-dimensional constraint on shadow radius and the dynamical stability analysis of gravitational perturbations demonstrates excellent agreement between these independent approaches. The conclusion of this cross-analysis further substantiates the accuracy of the high-dimensional shadow constraint formula proposed in this work, and we argue that this formula may serve as a universal tool for constraining the characteristic parameters of other high-dimensional spherically symmetric black hole solutions.
@@ II.1 The action and the field equations
$\mathcal{S}=\frac{1}{2}\int_{\mathcal{M}}\mathrm{d}^{n}x\sqrt{-g}\left[\mathcal{R}-\frac{(n-2)(n-1)}{3}\Lambda+{\alpha}_{2}\mathcal{L}_{\text{GB}}+\mathcal{L}_{\text{PYM}}\right],$ (1)
where $8\pi\mathscr{G}=1$, the parameter ${\alpha}_{2}$ is the Gauss-Bonnet coupling constant Cai (2002), and the Gauss-Bonnet Lagrangian $\mathcal{L}_{\text{GB}}$ is given as
$\mathcal{L}_{\text{GB}}=\mathcal{R}_{\gamma\zeta\varpi\delta}\mathcal{R}^{\gamma\zeta\varpi\delta}-4\mathcal{R}_{\gamma\zeta}\mathcal{R}^{\gamma\zeta}+\mathcal{R}^{2}.$ (2)
Here, the Lagrangian density of the power-Yang-Mills field Mazharimousavi and Halilsoy (2009); Ali and Saifullah (2021) is given by
$\mathcal{L}_{\text{PYM}}=-\left(\mathcal{F}\right)^{q},$ (3)
where the associated Yang-Mills invariant $\mathcal{F}$ is represented as
@@
where $\mathbf{A}^{(a)}$ or $A_{\zeta}{}^{(a)}$ are the Yang-Mills potentials of the gauge group SO(n1)SO(n-1), $\sigma$ is the coupling constant and $C_{(b)(c)}^{(a)}$ are interpreted as the structure constants of $\frac{(n-1)(n-2)}{2}$-parameter Lie group. Note that in the construction of high-dimensional spherically symmetric analytical solutions, the gauge group is generally chosen as SO(n1)SO(n-1) to ensure compatibility with the space-time symmetry. The field equation of the space-time metric $g_{\mu\nu}$ and the gauge potential $A_{\zeta}{}^{(a)}$ after the variation of action can be explicitly expressed as
@@ II.2 The high-dimensional black hole solution
The metric for high-dimensional ($n\geq 5$) static spherically symmetric is ansatz Mazharimousavi and Halilsoy (2009); Ali and Saifullah (2021)
@@
where $\mathfrak{h}_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ represents the line element of a (n2)(n-2)-dimensional hypersurface with the volume $\Sigma_{(n-2)}$. The space-time is higher dimensional and isotropic, characterized by a spherical horizon topology $\mathbb{S}^{n-2}$. The solution for a EPYMGB black hole in asymptotically flat space-time can be written as follows:
@@
This metric is the generalization of Schwarzschild space-time in higher dimensions cases, which is proved to be Ricci-flat. When considering an uncharged black hole, it can be observed that the EPYMGB solution (18) can be transformed into the Schwarzschild-Tangherlini solution (19) by applying the limit condition ${\alpha}_{2}\rightarrow 0$. When no limit condition is imposed on ${\alpha}_{2}$, the uncharged EPYMGB solution (18) is the EGB solution Cai (2002) in asymptotically flat space-time, which is
@@
When the limit ${\alpha}_{2}\rightarrow 0$ is set and $q=1$ is selected, the EPYMGB solution (18) will be reduced to the EYM solution Guo and Miao (2020) as follows:
@@
In order for the EPYMGB black hole solution to have at least one event horizon, that is, to ensure that the black hole does not expose a naked singularity, the values of the parameters ${\alpha}_{2}$ and $\mathcal{Q}$ are restricted to a fixed region, as shown in Figures 2 and 3. It is worth noting that in the process of numerical calculation of this fixed region, the specific conditions of the metric (16) should also be considered as follows:
@@ III.2 Standardization of shadow constraints
$\theta_{\mathrm{g}}=\frac{GM}{c^{2}D},$ (74)
where $G$ represents the gravitational constant term, $M$ is the black hole mass inferred from the EHT and $D$ is the observed distance between the black hole and the observer. Next, the fractional deviation of the angular gravitational radius is introduced as follows:
@@
where $\theta_{\mathrm{g}}$ and $\theta_{\mathrm{dyn}}$ are used to represent the real measurement and the inference of the stellar-dynamics of the angular gravitational radius, respectively. Although the definition $\delta={\theta_{\mathrm{g}}}/\theta_{\mathrm{dyn}}-1$ is referenced in some related papers, the two definitions are numerically similar; To maintain the clarity of the derivation, we adopt the expression $\delta=\theta_{\mathrm{dyn}}/{\theta_{\mathrm{g}}}-1$. Here, parameters $\theta_{\mathrm{g}}$ and $\theta_{\mathrm{dyn}}$ correspond to the measured value and estimated value, respectively. Thus, $\theta_{\mathrm{dyn}}$ can be written as
$\theta_{\mathrm{dyn}}=\frac{GM_{\mathrm{dyn}}}{c^{2}D_{\mathrm{dyn}}}.$ (76)
@@ III.2.2 High-dimensional space-time
Next, we define the angular gravitational radius in higher-dimensional space-time, denoted by
$\theta_{\mathrm{g}}=\frac{1}{D}\left(\frac{\mathscr{G}M}{c^{2}}\right)^{\frac{1}{n-3}},$ (87)
which is constructed to ensure dimensional consistency in arbitrary space-time dimensions. We further introduce the function $\mathfrak{C}(n)$, defined as follows:
@@
$=\mathfrak{C}(n)\cdot\left[\frac{1}{D_{\mathrm{dyn}}}\left(\mathscr{G}M_{\mathrm{dyn}}\right)^{\frac{1}{n-3}}\right]$
$=\mathfrak{C}(n)\cdot\theta_{\mathrm{dyn}}.$
@@
$\delta=\frac{\left[\frac{\theta_{\mathrm{sh}}}{\mathfrak{C}(n)}\right]}{\theta_{\mathrm{g}}}-1,$ (90)
@@
$\theta_{\mathrm{sh}}=\mathfrak{C}(n)\left(\delta+1\right)\theta_{\mathrm{g}}.$ (91)
@@
$\theta_{\mathrm{sh}}=\frac{r_{\mathrm{sh}}}{D}=\frac{r_{\mathrm{sh}}}{\left[\frac{\left(\mathscr{G}M\right)^{\frac{1}{n-3}}}{\theta_{\mathrm{g}}}\right]}=\mathfrak{C}(n)\left(\delta+1\right)\theta_{\mathrm{g}},$ (92)
@@
$r_{\mathrm{sh}}=\mathfrak{C}(n)\left(\delta+1\right)\left(\mathscr{G}M\right)^{\frac{1}{n-3}}.$ (93)
As defined above, it follows clearly that $M_{\mathrm{dyn}}\equiv\mathsf{M}$. In practical calculations, we adopt a system of units in which the ADM mass $\mathsf{M}$, the mass M{M}, and the high-dimensional gravitational constant $\mathscr{G}$ are set to unity, i.e., $\mathsf{M}=M=\mathscr{G}=1$. It should be emphasized that this choice does not imply that the physical value of $\mathscr{G}$ is identical across different space-time dimensions; rather, the gravitational constant in each dimension is simply used as a reference unit. Consequently, as long as one is not concerned with physical comparisons between different dimensions or with dimensional effects associated with the gravitational constant, this choice of units is entirely justified.
@@
The shadow radius constraint boundaries $r_{\mathrm{sh}}$ in high-dimensional space-time exhibit explicit dimensional dependence, with distinct parameter bounds emerging for each space-time dimension $n$. Additionally, an intriguing phenomenon is that for the high-dimensional constraint equation, the contribution of $\mathscr{G}$ to $r_{\mathrm{sh}}$ is affected by the dimension $n$ when $\mathscr{G}\neq 1$. It can be seen that Eqs. (93) and (84) are equivalent when the space-time dimension is set to four.
The confidence interval for the fractional deviation parameter $\delta$ is obtained here using the following measurement method. For $\mathrm{Sgr~A}^{\star}$, $\delta$ is computed relative to the expected shadow size from monitoring stellar orbits Akiyama and others (2022a); For $\mathrm{M87}^{\star}$, $\delta$ is computed for the expected shadow both from stellar dynamics and from gas dynamics Akiyama and others (2019b). Although the methods used to obtain the $\delta$ indirectly imply a gravitational framework governed by general relativity, the physical definition of the parameter $\delta$ does not mandate that the observed black hole space-time must be described and interpreted using general relativity. Therefore, it can be considered that the parameter $\delta$ itself does not depend on a specific theoretical model. This is reflected in the fact that the EHT employs certain parameterized test metrics as well as modified metrics to examine whether the black hole geometry deviates from the Schwarzschild (or Kerr) metric Akiyama and others (2022b); Psaltis and others (2020); Kocherlakota and others (2021). If $\delta\simeq 0$ is theoretically satisfied, then it can be definitively assumed that the space-time geometry of the black hole is described by the corresponding Schwarzschild (or Kerr) solution of general relativity. Therefore, the effective range of $\delta$ obtained with the current observational accuracy still allows for the possibility that other metrics may exist within reasonable ranges of their characteristic parameters Khodadi and Saridakis (2021); Uniyal et al. (2023); Xavier et al. (2023); Pantig et al. (2022, 2023).
Furthermore, when considering black hole space-times with extra dimensions, if the Schwarzschild (or Kerr) metric adopted by the EHT is used as the default benchmark for modeling, the corresponding parameter $\delta$ should systematically deviate from $\delta\simeq 0$. Here we speculate that the confidence interval of $\delta$ may already include the deviation induced by the presence of extra dimensions, and that the higher-dimensional deviation to the shadow remains smaller than the current observational uncertainty. Based on this conjecture, we proceed to investigate two classes of high-dimensional black hole space-times: black string space-times and general high-dimensional spherically symmetric black hole space-times. For black string space-times, their geometric structure can be effectively regarded as a four-dimensional Schwarzschild space-time once the extra dimensions are compactified. Accordingly, we still adopt the Schwarzschild (or Kerr) metric as the benchmark for modeling. For the space-time of a general spherically symmetric high-dimensional black hole, since the high-dimensional shadow is projected onto a two-dimensional $(\mathrm{X},\mathrm{Y})$ plane, the influence of higher dimensions on the shadow scale is encoded in the space-time metric. In other words, the metric that serves as the baseline for constraint modeling should be highly correlated with the metric used to compute the high-dimensional shadow radius. Therefore, when discussing the EPYMGB metric, which deviates slightly from the Schwarzschild-Tangherlini metric, we adopt the Schwarzschild-Tangherlini metric modeling as the benchmark.
@@ IV.1.4 Spin-2 perturbations
During the calculations, it is important to emphasize that gravitational perturbations in the EPYMGB space-time are extremely complex, rendering the derivation of the corresponding master equations technically challenging. Moreover, the physically admissible ranges of the parameters ${\alpha}_{2}$ and $\mathcal{Q}$ are considerably restricted, which leads to only a slight deviation between the EPYMGB solution and the Schwarzschild-Tangherlini solution when considering cases with small parameter values. Therefore, this approximation is well justified when employing the gravitational perturbation equations derived from the higher-dimensional vacuum Einstein field equations, although it may omit certain features of the full gravitational spectrum.
@@ V.2 High-dimensional constraint results of shadow radius
Firstly, we apply the high-dimensional shadow radius formula (72) constructed in this paper to investigate the EPYMGB black hole solutions, and we present the corresponding numerical result of $R_{s}$ within the allowed parameter space $\left({\alpha}_{2},\mathcal{Q}\right)$. Subsequently, we utilize the high-dimensional shadow constraint formula (93) to provide constraints on the previously obtained values of $R_{s}$ based on observational data from the $\mathrm{M87}^{\star}$ and $\mathrm{Sgr~A}^{\star}$, respectively. The results are shown in Figures 4, 5, 6 and 7. Among these results, we set the ADM mass of the black hole and the four-dimensional gravitational constant as $\mathsf{M}=M=1$ and $\mathscr{G}=1$, respectively.